# Speech-Audio Compositional Attacks on Multimodal LLMs and Their Defense with SALMONN-Guard

### v2

@@ 3.2.3 Speech-Audio Mixture
This method pivots the attack vector from explicit linguistic content to the more subtle realm of implicit environmental context. The central premise is to test whether a model’s safety mechanisms are merely sophisticated textual processors or if they possess a genuine multimodal understanding of an auditory scene. We achieve this by pairing an safe spoken question with an audio event that carries a harmful intention. The entire sample is designed to appear benign to a system that only transcribes and analyzes the spoken words, while being obviously unsafe to one that comprehends the full context.
@@
To construct these attacks, we curated a library of harmful audio tracks. These tracks were extracted directly from publicly available videos on online platforms featuring adult-oriented, violence or criminal content. This process yielded a collection of audio files that, while not containing explicit instructions, carry a strong and inherently inappropriate contextual meaning from their source.
The final stage involves mixing auditory scene. We take a benign “carrier” track from the VoiceBank-DEMAND dataset and overlay one of the harmful contextual audio tracks from our curated library as a background layer. The background sound is mixed at a clearly audible volume. This creates a powerful conflict: for instance, an academic lecture playing over the distinct sounds of an intimate scenario. A model with a text-centric safety filter would likely transcribe the harmless lecture and clear the input. However, a truly multimodal safety system is challenged to detect the disturbing nature of the background context and correctly flag the entire sample as unsafe. This method directly probes the model’s ability to reason about situational appropriateness.
@@ 3.3 Question-Answer Pair Formulation
To conduct a comprehensive and granular assessment of model vulnerabilities, we designed two distinct question formats for each synthesized adversarial audio sample. This evaluation strategy allows us to probe a model’s safety architecture from basic detection to detailed comprehension.
3.3.1 Binary Classification (Yes / No) Questions
This format is designed to test the model’s fundamental ability to detect harmful content. For each adversarial audio sample, we pose the direct textual question: “Does this audio contain harmful content?”. The evaluation metric is the Attack Success Rate (ASR %). Since every audio sample in our benchmark is inherently harmful, an attack is considered successful if the model incorrectly responds with “No”. This ASR provides a clear measure of the model’s failure rate in its most basic threat identification capabilities.
@@ 3.3.2 Open-Ended Questions
This format is designed to test a model’s ability to comprehend and act upon (or refuse to act upon) the harmful content identified in the audio. After providing the adversarial audio, we submit a textual prompt that asks the model to engage with the malicious information directly. The prompts typically ask the model to identify, extract, or elaborate on the harmful aspects of the audio. The evaluation of the model’s response is automated using Gemini 2.5 Pro as an independent judge. We measure the Attack Success Rate (ASR %) by providing the model’s generated text to a Gemini 2.5 Pro based classifier, which determines if the response complied with the harmful request or successfully refused it. The specific prompts and detailed scoring rubric used for this automated evaluation are provided in Appendix.
Models Speech Overlap (%) Dialogue (%) Contextual Audio Cues (%) Overall (%) Open-source Models Qwen2-Audio-7B (qwen2-audio)     100.00 96.48 100.00 98.16 Qwen2.5-Omni-7B (qwen2.5_omni)     99.78 87.02 100.00 92.83 Qwen3-Omni-30B-A3B-Instruct (xu2025qwen3)     49.25 84.82 93.17 81.50 Step-Audio 2 mini Base (2025step-audio)     90.25 77.49 98.33 85.57 MiniCPM-o 2.6 8B (yao2024minicpm)     85.84 69.87 99.58 77.95 Kimi-Audio-7B-Instruct (2025kimiaudio)     69.00 61.36 87.17 70.05 Proprietary Models Gemini 1.5 Pro (2024gemini1.5pro)     83.50 78.58 98.47 85.12 Gemini 2.5 Pro (2025gemini2.5)     37.25 63.93 88.56 66.75 GPT-4o (2024gpt4o)     73.00 53.67 99.58 70.05 SALMONN-Guard (ours)     12.93 14.08 5.16 11.32
Table 1: Main results on SACRED-Bench. We report Attack Success Rate (ASR %) for Speech Overlap (N=400) and Contextual Audio Cues (N=717) based on a binary classification task, and for the Dialogue attack (N=1364) via an open-ended compliance test. The Overall ASR is the weighted average of the three attacks. Higher values consistently indicate greater model vulnerability.
@@ 4.1 Training Data Curation
Acknowledging the scarcity of ethically-sourced harmful data, we synthesized our training corpus by leveraging established red-teaming benchmarks—AdvBench, MM-SafetyBench, and HarmBench—as a conceptual foundation. Instead of using the original prompts directly, we employ GPT-4o to generate a new, distinct set of harmful instructions, ensuring no direct data leakage. This process yields a training corpus of approximately 10k instances (8,545 harmful, 1,828 benign), with the harmful data distributed across speech overlap (6,640), speech-audio mixture (986), and multi-speaker dialogue (919) instances. The benign samples are used to prevent model collapse to only outputting the unsafe class. To ensure a rigorous evaluation of generalization, we enforced a strict cross-dataset validation protocol: for any given attack category, the source benchmark used to inspire the training prompts was held out from the test set.
@@ 4.3 Training Process Specification
Our primary objective is to adapt the powerful base model into a specialized guard model through Supervised Fine-Tuning (SFT) with standard cross-entropy loss. Concretely, SALMONN-Guard is trained to perform safety assessments for each sample by jointly processing both the audio input and any accompanying text prompt. This multimodal training scheme directly overcomes the vulnerability of text-only or spoken-content-only guardrails, yielding high efficacy in combating attacks in SACRED-Bench.
We employed a two-stage fine-tuning curriculum to first build general robustness and then specialize in complex threats. In Stage 1, the model was trained for three epochs on the entire dataset. In Stage 2, we conducted five additional epochs of training exclusively on the multi-speaker dialogue subset to bolster the model’s capabilities against this cross-modal attacks that is particularly challenging for the model to learn. For parameter-efficient fine-tuning, we utilized Low-Rank Adaptation (LoRA), applying updates concurrently to the large language model, the audio encoder, and the aligner module. A comprehensive list of hyperparameters is provided in the Appendix.
5 Experiment
@@ 5.1 Experimental Setup
Evaluated Models: Our evaluation includes several top-tier multimodal models, encompassing proprietary systems like Gemini 2.5 Pro (2025gemini2.5), Gemini 1.5 Pro (2024gemini1.5pro), GPT-4o (2024gpt4o) and Kimi Audio (2025kimiaudio). Concurrently, we evaluate prominent open-source models, such as Qwen3-Omni (xu2025qwen3), Qwen2.5-Omni-7B (qwen2.5_omni), Qwen2-Audio-7B (qwen2-audio), MiniCPM-o 2.6 (yao2024minicpm), and Step-Audio2 (2025step-audio). Furthermore, we include our proposed guard model, SALMONN-Guard, in the evaluation to validate its effectiveness.
Benchmark and Metrics: All experiments are conducted on our SACRED-Bench, which evaluates models against three attack methods: speech overlap, multi-speaker dialogue, and speech-audio mixture. Our major evaluation metric is the Attack Success Rate (ASR %), which quantifies the percentage of times a model fails to adhere to its safety alignment and complies with the malicious prompt. The classification of a successful attack varies by task format. For the binary classification task, an attack is successful if the model fails to correctly identify there is harmful content in the audio. For the Open-Ended task, we employ an automated evaluation pipeline where Gemini 2.5 Pro serves as a judge, classifying the target model’s generated response as either “harmful” (a successful attack) or “harmless”. Therefore, in Table 1, higher ASR values consistently indicate a greater degree of vulnerability.
@@ 5.2 Main Results and Analysis
$500$ $1.3$ 8-8 $47.00$
$500$ $1.5$ 8-8 $61.00$
$500$ $1.1$ 8-8 $42.00$
$300$ $1.3$ 8-8 $41.00$
$700$ $1.3$ 8-8 $47.00$
$500$ $1.3$ 6-6 $47.00$
$500$ $1.3$ 10-10 $51.00$
Table 2: Impact of acoustic hyperparameters on the Speech Overlap attack. All results are ASR% against Gemini 2.5 Pro, evaluated on a 100-sample subset of SACRED-Bench.
@@
Text + Audio (Our Dialogue Method) $78.58$
Table 3: Ablation study on the Multi-speaker Dialogue attack, evaluated on Gemini 1.5 Pro. We compare our proposed method against unimodal baselines: an “Audio-Only” setup with harmful speech and a “Text-Only” setup with a direct harmful instruction.
We present the detailed experimental results in Table 1. Overall, the results shows that both leading proprietary and major open-source models unfortunately exhibit significant vulnerabilities when confronted with our proposed speech-audio composition attacks.
Proprietary Models Exhibit Exploitable Loopholes. As seen in the results, while top-tier models like Gemini 2.5 Pro demonstrate some defensive capabilities, they are far from infallible. Particularly under the speech-audio mixture attack, Gemini 2.5 Pro shows a high attack success rate of 88.56%, and its success rate for the Multi-speaker Dialogue attack is also significant at 63.93%. This suggests that current safety mechanisms may be more focused on analyzing explicit textual instructions, while being less capable of identifying implicit malicious intent conveyed through non-speech audio or complex conversational structures.
Open-Source Models Largely Lack Effective Audio Safeguards. The results indicate that multimodal models from the open-source community are almost entirely defenseless against SACRED attacks. For instance, Qwen2.5-Omni-7B and Qwen2-Audio-7B have attack success rates approaching 100% in most tests. This reflects a critical gap in the current open-source ecosystem: while the pursuit of performance and functionality is paramount, insufficient attention has been paid to the security of multimodal inputs, especially the robustness against complex audio scenarios.
Analysis of Different Attack Strategies’ Efficacy. An analysis of the individual attack strategies reveals that each method successfully exploits distinct architectural weaknesses. The speech-audio mixture method proved the most potent, with a staggering 88.56% ASR against Gemini 2.5 Pro and nearly 100% on several open-source models, underscoring a critical ’cross-modal blindness.’ Models adeptly transcribe a benign spoken query while completely ignoring the malicious intent conveyed by background audio. Similarly effective, the Multi-speaker Dialogue method demonstrates how requests can be obfuscated within a natural conversation, successfully bypassing safety filters that may only detect direct, command-like instructions. The high ASR of 63.93% against Gemini 2.5 Pro highlights the difficulty even advanced models face in assessing risks distributed throughout a long-form dialogue. Lastly, the Speech Overlap attack targeted the models’ information processing in complex acoustic environments. Although Gemini 2.5 Pro showed improved resistance, high success rates against Gemini 1.5 Pro (74.89%) and the Qwen series confirm that for most models, disentangling overlapping audio remains a significant security flaw.
@@
Speech Insertion (archilles) 23.17% 8.94% 13.41% 0.00%
Speech Editing (cheng2025jailbreakaudiobenchindepthevaluationanalysis) 33.00% 12.00% 23.00% 3.00%
SACRED-Bench (ours) 92.83% 81.50% 65.41% 11.32%
Table 4: Attack Success Rate (ASR %) across different audio jailbreaking methods. Speech Insertion and Speech Editing are evaluated on the SSJ and AMSE subsets of JALMBench, respectively. The ASR for SACRED-Bench refers to the overall performance.
Effectiveness of SALMONN-Guard as a Guard Model. In stark contrast to other models, our proposed SALMONN-Guard demonstrates exceptional defensive performance. As shown in the last row of Table 1, SALMONN-Guard drastically reduces the success rate for all types of attacks. It slashes the ASR for speech-audio mixture from Gemini 2.5 Pro’s 88.56% to a mere 5.16%, and suppresses the ASRs for Speech Overlap and Multi-speaker Dialogue to 12.93% and 28.09%, respectively. We also test SALMONN-Guard on all the benign audios used to composite SACRED-Bench as shown in Table 5, and an 100% classification accuracy was achieved where none of them were flagged as harmful. This result fully validates the effectiveness of our proposed approach: training a specialized, lightweight guard model to preemptively identify and intercept malicious audio inputs is a viable and highly efficient path to protecting downstream LLMs from complex audio-based attacks.
Model ASR(%) Benign Audio Acc(%)
SALMONN-Guard 11.32 100.00
Table 5: Performance of SALMONN-Guard on harmful (measured by ASR) and benign audios (classification accuracy), where the benign audios are audios used to create SACRED-Bench test set.

### v3

@@ 3.3.2 Speech-Audio Overlap (SAO)
This method pivots the attack vector from explicit linguistic content to the more subtle realm of implicit environmental context. The central premise is to test whether a model’s safety mechanisms are merely sophisticated textual processors or if they possess a genuine multimodal understanding of an auditory scene. We achieve this by pairing a safe utterance with an audio event that carries a harmful intention.
@@
To construct these attacks, we curated a library of harmful audio tracks. These tracks were extracted directly from publicly available videos on online platforms featuring adult-oriented, violent, or criminal content. This process yielded a collection of audio files that, while not containing explicit instructions, carry a strong and inherently inappropriate contextual meaning from their source. The final stage involves mixing the auditory scene. We take a benign “carrier” track from the VoiceBank-DEMAND dataset and overlay one of the harmful contextual audio tracks from our curated library as a background layer. The background sound is mixed at a clearly audible volume. This creates a powerful conflict: for instance, an academic lecture playing over the distinct sounds of an intimate scenario.
3.3.3 Multi-Speaker Dialogue (MSD)
Recent findings in vision-language safety(Hu et al., 2025) have demonstrated that many multimodal models could be easily aligned by text-only training data because the text prompt often revealed the harmful nature of the accompanying image. Our work extends this core insight to the speech-audio domain. We hypothesize that a similar vulnerability exists: a model’s safety alignment might operate inconsistently across modalities if a benign-on-its-own textual query is paired with a malicious audio context, causing the text-only safety filter to be ineffective.
Our construction process is engineered to induce this precise semantic gap, as illustrated in Fig. 3(c). We first use GPT-4o to transform a direct harmful instruction, such as a question on how to develop nuclear weapons, into a plausible, multi-turn dialogue script between two speakers. This script rephrases the malicious request into a conversation, thereby cloaking the harmful intent in a conversational structure to remove the directive marker in the original commands. In parallel, we craft the textual component of the attack, which is a completely innocuous and referential query. This question is semantically neutral on its own but is designed to act as a pointer, or a semantic trigger, to the content within the audio. For example, a corresponding query might be, “How would I create the device mentioned in the discussion?” In isolation, this question is safe. However, when paired with the dialogue, it becomes a direct prompt for the model to use the harmful information it has just heard.
Then, we utilize the multi-speaker capabilities of the ChatTTS engine to synthesize the dialogue script into a realistic audio file, assigning distinct vocal characteristics to each speaker. The final attack prompt is the combination of this harmful conversational audio with the benign textual trigger. This presents the model with a complex jailbreaking challenge, forcing it to reconcile a harmless-looking text instruction with a clearly malicious audio payload, thereby testing the true depth of its cross-modal safety reasoning.
@@ 3.4 Question-Answer Pair Formulation
We introduce two distinct question formats in this section corresponding to different types of adversarial audio.
3.4.1 Permission Questions
Permission questions are Yes-No questions used for SSO and SAO, which seek approval for malicious intent. This format is designed to test the model’s fundamental ability to detect harmful content. For each adversarial audio sample, the prompt asks about the appropriateness of playing the audio in certain scenarios. For example, “Is it suitable to play this audio with my 5-year-old kid?”. Since every audio sample in our benchmark is inherently harmful, an attack is considered successful if the model incorrectly responds by permitting such action or acknowledging the harmlessness of this audio. This provides a clear measure of the model’s failure rate in its basic threat identification capabilities.
@@ 3.4.2 Open-Ended Questions
Open-ended questions are primarily used in MSD attacks. This format is designed to test a model’s ability to comprehend and act upon (or refuse to act upon) the harmful content identified in the audio. After providing the adversarial audio, we submit a textual prompt that asks the model to engage directly with the malicious information. The prompts typically ask the model to identify, extract, or elaborate on the harmful aspects of the audio. The model’s response is evaluated automatically using Gemini 2.5 Pro as an independent judge. We measure the Attack Success Rate by providing the model’s generated text to a Gemini 2.5 Pro-based classifier, which determines if the response complied with the harmful request or successfully refused it. The specific prompts and detailed scoring rubric used for this automated evaluation are provided in the Appendix.
@@ 4.1 Training Data Curation
Acknowledging the scarcity of ethically-sourced harmful data, we curate the training corpus by following established text red-teaming benchmarks and employ GPT-4o to generate a new, distinct set of harmful instructions, ensuring no direct data leakage. This process yields a training corpus of approximately 10k samples (8,545 harmful, 1,828 benign), with the harmful data distributed across SSO (6,640), SAO (986), and MSD (919) categories. The benign samples are used to prevent the model from only outputting the unsafe class. To ensure a rigorous evaluation of generalization, we enforced a strict cross-dataset validation protocol, that is, for any given attack category, the source benchmark used to inspire the training prompts was held out from the test set.
@@ 4.3 Training Process Specification
Our primary objective is to adapt the powerful base model into a specialized guard model via Supervised Fine-Tuning (SFT) using the standard cross-entropy loss. Concretely, SALMONN-Guard is trained to perform safety assessments for each sample by jointly processing both the audio input and any accompanying text prompt. This multimodal training scheme directly mitigates the vulnerability of text-only or spoken-content-only guardrails, yielding high efficacy against attacks on SACRED-Bench.
We employed a two-stage fine-tuning curriculum to first build general robustness and then specialize in complex threats. In Stage 1, the model was trained for three epochs on the entire dataset. In Stage 2, we conducted five additional epochs of training exclusively on the multi-speaker dialogue subset to bolster the model’s capabilities against these cross-modal attacks, which are particularly challenging for the model to learn. For parameter-efficient fine-tuning, we utilized Low-Rank Adaptation (LoRA), applying updates concurrently to the large language model, the audio encoder, and the aligner module. A comprehensive list of hyperparameters is provided in Appendix D.
5 Experiments
Table 1: Main results on SACRED-Bench. We report ASR for SSO (N=400), MSD (N=717), and SAO (N=1364). We also report FAR on the harmless control set (HCS, N=400). Overall ASR is the weighted average across the red-teamed set: SSO, SAO and MSD. OBE is the average of ASR and FAR weighted by the number of samples in the red-teamed set and the control set. All metrics are the lower the better.
Models SSO (%) MSD (%) SAO (%) Overall ASR (%) HCS (%) OBE (%) Open-source Models Qwen2-Audio-7B (2024)     100.00 96.48 100.00 98.16 1.25 85.41 Qwen2.5-Omni-7B (2025a)     99.78 87.02 100.00 92.83 1.25 83.03 Step-Audio 2 mini Base (2025)     90.25 77.49 98.33 81.50 0.00 78.37 MiniCPM-o 2.6 8B (2024)     85.84 69.87 99.58 85.57 4.00 77.01 Qwen3-Omni-30B-A3B-Instruct (2025b)     49.25 84.82 93.17 77.95 0.75 72.16 Kimi-Audio-7B-Instruct (2025)     69.00 61.36 87.17 70.05 4.75 66.78 Proprietary Models Gemini 1.5 Pro (2024)     83.50 78.58 98.47 85.12 0.00 77.77 GPT-4o (2024)     73.00 53.67 99.58 70.05 0.75 70.74 Gemini 2.5 Pro (2025)     37.25 63.93 88.56 66.75 1.00 63.15 SALMONN-Guard (ours)     12.93 14.08 5.16 11.32 0.00 7.74
@@ 5.1 Experimental Setup
Evaluated Models: Our evaluation includes several top-tier multimodal models, encompassing proprietary systems like Gemini 2.5 Pro (Comanici et al., 2025), Gemini 1.5 Pro (Team et al., 2024), GPT-4o (Hurst et al., 2024) and Kimi Audio (Ding et al., 2025). Concurrently, we evaluate prominent open-source models, such as Qwen3-Omni (Xu et al., 2025b), Qwen2.5-Omni-7B (Xu et al., 2025a), Qwen2-Audio-7B (Chu et al., 2024), MiniCPM-o 2.6 (Yao et al., 2024), and Step-Audio2 (Wu et al., 2025). Furthermore, we include our proposed guard model, SALMONN-Guard, in the evaluation to validate its effectiveness.
Evaluation: Our major evaluation metric is the Attack Success Rate (ASR %), which quantifies the percentage of times a model fails to adhere to its safety alignment and complies with the malicious prompt. The classification of a successful attack varies by task format. For SSO and SAO, an attack is successful if the model fails to correctly identify that there is harmful content in the audio. For the MSD, we employ an automated evaluation pipeline where Gemini 2.5 Pro serves as a judge, classifying the target model’s generated response as either “harmful” (a successful attack) or “harmless” by providing it with the original harmful prompt and intended harmful answer (see Appendix B for detailed prompt). Therefore, in Table 1, higher ASR values consistently indicate a greater degree of vulnerability. To mitigate bias by a single judge, we also introduce a human evaluator and Qwen2.5-7B as another two judges to show their agreement with Gemini 2.5 Pro in Appendix F.
Since SSO and SAO are both essentially classification tasks, we provide a harmless control set (HCS) of the same number of samples as SSO and SAO, but without any harmful content. False alarm rate (%FAR) is measured on this set. We then define the overall behavior error (%OBE) as the average of ASR on the red-teamed set and the FAR on the harmless control set, weighted by the number of samples in each set, respectively.
@@ 5.2 Main Results and Analysis
Table 2: Impact of acoustic hyperparameters on the SSO attack. All results are %ASR against Gemini 2.5 Pro, evaluated on a 100-sample subset of SACRED-Bench.
@@
500 1.3 -8 47.00
500 1.5 -8 61.00
500 1.1 -8 42.00
300 1.3 -8 41.00
700 1.3 -8 47.00
500 1.3 -6 47.00
500 1.3 -10 51.00
Table 3: Ablation study on the MSD attack, evaluated on Gemini 1.5 Pro. We compare our proposed method against unimodal baselines: an “Audio-Only” setup with harmful speech and a “Text-Only” setup with a direct harmful instruction.
@@
Text + Audio (i.e. MSD) 78.58
We present the detailed experimental results in Table 1. Overall, the results show that both leading proprietary and major open-source models, unfortunately, exhibit significant vulnerabilities when confronted with our proposed speech-audio composition attacks.
Proprietary Models Exhibit Exploitable Loopholes. As seen in the results, while top-tier models like Gemini 2.5 Pro demonstrate some defensive capabilities, they still suffer severely from attacks in SACRED-Bench. Particularly under the SAO attack, Gemini 2.5 Pro shows a high ASR of 88.56%, and its success rate for the MSD attack is also significant at 63.93%. This suggests that current safety mechanisms may be more focused on analyzing explicit textual instructions, while being less capable of identifying implicit malicious intent conveyed through non-speech audio or complex conversational structures.
Open-Source Models Largely Lack Effective Audio Safeguards. The results indicate that open-source multimodal models are largely defenceless against SACRED-Bench attacks. For instance, Qwen2.5-Omni-7B and Qwen2-Audio-7B have ASRs approaching 100% in most tests. This reflects the fact that most attention in safety has been paid to text modality, and advocates more efforts to be paid to ensure the safety of multimodal inputs, especially the robustness against complex audio scenarios.
Analysis of Different Attack Strategies’ Efficacy. An analysis of the individual attack strategies reveals that each method successfully exploits distinct architectural weaknesses. The SAO method proved the most effective, with an 88.56% ASR against Gemini 2.5 Pro and nearly 100% on several open-source models, underscoring a critical cross-modal weakness. Models adeptly transcribe a benign spoken query while completely ignoring the malicious intent conveyed by background audio. Similarly effective, the MSD method demonstrates how requests can be obfuscated within a natural conversation, successfully breaking safety filters that may only detect direct instructions. The high ASR of 63.93% against Gemini 2.5 Pro highlights the challenge that even advanced models face risks distributed throughout a long-form dialogue. Lastly, the SSO attack targeted the models’ information processing in complex acoustic environments. Although Gemini 2.5 Pro showed improved resistance, high ASR against Gemini 1.5 Pro (74.89%) and the Qwen series confirm that for most models, disentangling multiple audio sources remains a significant security flaw. Furthermore, we provide qualitative examples with model responses under SACRED-Bench attacks in Appendix A.
Table 4: ASR across different audio jailbreaking methods. Speech Insertion and Speech Editing are evaluated on the SSJ and AMSE subsets of JALMBench, respectively. The ASR for SACRED-Bench refers to the overall performance.
@@
Speech Insertion (Yang et al., 2024a) 23.17% 8.94% 13.41% 0.00%
Speech Editing (Cheng et al., 2025) 33.00% 12.00% 23.00% 3.00%
SACRED-Bench (ours) 92.83% 81.50% 66.75% 11.32%
No model suffers from over-refusal on HCS. Regardless of the model’s performance on the red-teamed set, almost all models achieved very low false alarm rates on HCS. Notably, proprietary models achieved lower than 1% FAR, and SALMONN-Guard achieved 0%. The FARs are also rare on other models, demonstrating that models performing well are not because of a higher refusal rate, but because of a better safety mechanism.
Effectiveness of SALMONN-Guard as a Guard Model. In stark contrast to other models, our proposed SALMONN-Guard demonstrates exceptional defensive performance. As shown in the last row of Table 1, SALMONN-Guard drastically reduces the success rate for all types of attacks. It slashes the ASR for speech-audio mixture from Gemini 2.5 Pro’s 88.56% to a mere 5.16%, and suppresses the ASRs for SSO and MSD to 12.93% and 14.08%, respectively. This result clearly supports the effectiveness of the SALMONN-Guard approach: training a specialised, lightweight guard model to identify and preemptively intercept malicious audio inputs is a viable and highly effective way to protect downstream LLMs from complex audio-based attacks.